A stone is projected vertically upward with a viscosity of 100ms-1  at an angle of 30⁰ to the radius 1mm is extended by 25mm on the application of a force of 104N. Calculate (i) Young modulus (ii) strain energy per unit volume.


1
Expert's answer
2021-03-08T07:32:06-0500

i)


"Y=\frac{\sigma}{\epsilon}\\Y=\frac{Fl}{\pi r^2\Delta l}
\\Y=\frac{10^4(1)}{\pi (0.001)^2(0.025)}=1.3\cdot10^{11}Pa"

ii)


"E=\frac{10^4(0.025)}{2\pi (0.001)^2(1)}=4.0\cdot10^7\frac{J}{m^3}"